OpenStudy (anonymous):

A car starts from rest and accelerates at 0.3 m/s^2. What is the speed of the car after it has traveled 25m?

OpenStudy (anonymous):

Use the following relations firstly to find time t and secondly put it in the later equation to find the desired result. x=1/2at^2 v=at

OpenStudy (anonymous):

25=1/2*0.3sqrt -->t=.... now put the value of t in v=0.3 t ....